OpenStudy (anonymous):

plzzzzzzzz helpppppppp!!! 3+sqrt{z-9}=sqrt{z+6}

OpenStudy (anonymous):

square both sides to get \[3+\sqrt{z-9}=\sqrt{z+6}\] \[3^2+2*3*\sqrt{z-9}+z-9=z+6\] \[6*\sqrt{z-9}+z=z+6\] \[6*\sqrt{z-9}=6\] \[\sqrt{z-9}=1\] \[z-9=1 so z=10\]

OpenStudy (anonymous):

so z= 10